### v1

# Submitted to Astronomy Letters THE ERUPTIVE YOUNG STAR IRAS 21204+4913

Abstract
The results of photometric, polarimetric, and spectroscopic observations are presented for the young star IRAS 21204+4913, whose visible brightness has increased by since October 2025. The star’s absorption spectrum in the m range resembles those of A–F giants and supergiants, but it also exhibits molecular TiO bands. The brightening was accompanied by a significant increase in the degree of polarization of the stellar radiation (to in the -band), likely due to scattering by dust in an expanding circumstellar shell. The P Cygni profile of the H line implies a dusty wind velocity of km s-1. We believe that the outburst of IRAS 21204+4913 is caused by an increase in the accretion rate of protoplanetary disk’s matter onto the young star with a mass of M⊙ to M⊙ yr-1. Furthermore, IRAS 21204+4913 displays several unusual features: the dependence of the width and radial velocity of absorption lines on the excitation potential, emission in the TiO molecular bands, and a comparably bright outburst that occured in 1948. Several T Tauri stars and a group of Herbig-Haro objects are found in the vicinity of the star.
1 Introduction
During the transformation of a protostellar cloud into a T Tauri star – a process lasting Myr – some young stellar objects (YSOs) undergo dramatic brightening events that can persist from several months to several decades (Herbig, 1977). It is generally accepted that outbursts of this kind are caused by a strong increase in the accretion rate of protoplanetary disk matter onto the forming star (Hartmann and Kenyon, 1996, Hartmann et al., 2016, Audard et al., 2014). In extreme cases, can exceed M⊙ yr-1, which is thousands of times more than typical observed in classical T Tauri stars (CTTSs). If such events recur frequently during star formation, they may significantly influence the final stellar mass and other fundamental parameters of the star – see Baraffe et al. (2017), Fischer et al. (2023) and references therein. A strong, albeit relatively short-term, increase in the luminosity of YSOs undoubtedly influences the evolution of their protoplanetary disks, and therefore the formation of planetary systems – see, e.g., Stamatellos et al. (2012), Topchieva et al. (2025).
Various mechanisms have been proposed to explain the origin of powerful episodic accretion. Without aiming for completeness (see e.g. Fischer et al. 2023), we mention only a few: thermal, gravitational, and/or magnetorotational disk instabilities (Zhu et al., 2010, Bae et al., 2014, Meyer et al., 2017, Kadam et al., 2020); interactions between the disk and a planet or stellar companion (Borchert et al., 2022, Nayakshin et al., 2023); an infall of gas cloud onto the disk (Demidova and Grinin, 2023).
Depending on the outburst amplitude, duration, and spectral characteristics, Herbig (1977) divided eruptive YSOs into two classes: FU Orionis-type objects – FUors111The term fuor was introduced by V. A. Ambartsumyan (1971). and EX Lupi-type objects – EXors (Herbig, 1989). Herbig classified FUors as objects whose visible brightness increased by over several months and then gradually decreased over several decades, and whose spectra were almost entirely absorption-dominated throughout the outburst. EXors, in contrast, exhibit recurrent outbursts – sometimes reaching FUor-like amplitudes but lasting only weeks to a year – during which their spectra are dominated by emission lines.
Over the past decade, the advent of numerous monitoring systems has led to the discovery of a large number of eruptive YSOs. A team of researchers Contreras Peña,Carlos et al. (2025) compiled the OYCAT (Outbursting YSOs Catalogue)222http://starformation.synology.me:5002/OYCAT/main.html, which currently contains basic information on 174 objects. According to the established criteria (Contreras Peña,Carlos et al., 2025, table 1), the objects in the catalogue are divided into several groups based on the photometric and spectral characteristics of their outbursts. Without going into detail, we note that the authors confidently classified as FUors and EXors fewer than 40 % of all objects, whereas nearly half of the YSOs exhibited outburst parameters intermediate between those of FUors and EXors.
The diversity of observed manifestations of episodic accretion can be attributed either to different underlying physical mechanisms or to the combined action of multiple mechanisms – see, for example, Skliarevskii and Vorobyov (2023). Acquiring the most comprehensive possible information on each eruptive YSO will help elucidate the general physics of the phenomenon. In this context, we became interested in the report by S. Kaneko333https://www.astroarts.co.jp/photo-gallery/photo/129147, who discovered on November 25, 2025, the optical transient TCP J21220926+4926242 (hereafter J2122), whose brightness in the optical band approximately doubled over the following week.
Subsequently, it became clear that the outburst occurred no later than October 25, 2025, when the transient was first detected by the ASAS-SN survey (Shappee et al., 2014) at a magnitude of (Kochkina et al., 2025). By late November, when the brightness of J2122 exceeded , Kochkina et al. (2025) carried out the first spectroscopic observations of the transient in the wavelength range of m and found that its spectrum resembled the absorption-dominated spectra typical of FUor-like objects: the only emission feature detected in the spectrum of J2122 was the [S ii] 6731 Å line.
Grosso et al. (2025) identified the transient J2122 with the IRAS source 21204+4913 (Neugebauer et al., 1984), which, according to Marton et al. (2016), is a class I/II YSO – i.e., it is in the transitional evolutionary stage between a protostar and a pre-main-sequence star (Adams et al., 1987, Lada, 2005). IRAS 21204+4913 lies on the southeastern edge of the relatively compact (18 square arcminutes) dark nebula D 2944 (Dobashi, 2011). Grosso et al. (2025) determined a distance to the source of pc, and Stecklum (2025) reported a line-of-sight extinction of
We began observing IRAS 21204+4913 in early December 2025, and in this paper we present the results of our research. First, we describe our observational data and the results derived from them; then, we examine what our data reveal about the nature of IRAS 21204+4913. The conclusions and main findings of this work are summarized in the Conclusion.
2 Observations
To construct the historical light curve of IRAS 21204+4913, we selected approximately 100 photographic plates from the collection of the Sternberg Astronomical Institute of Lomonosov Moscow State University (SAI MSU) that included the star. These plates were obtained between May 10, 1899, and November 7, 1975, in a photometric system close to the Johnson band. The limiting magnitude of the plates varied from to , depending on the instrument used and the observing conditions.
The observations presented below were carried out in 2025–2026 with the instruments of the Caucasian Mountain Observatory (CMO) of SAI MSU (Shatsky et al., 2020). Optical photometry of IRAS 21204+4913 was obtained with the 60-cm RC600 telescope equipped with a CCD camera and a set of filters in the Bessell-Cousins photometric system (Berdnikov et al., 2020). The magnitudes of the comparison stars were calibrated against the standard field GD391 (Clem and Landolt, 2016). The results of the optical photometry are presented in Table A1.
Photometric observations of the star in the near-infrared (NIR) were carried out in the bands of the MKO-NIR photometric system using the ASTRONIRCAM infrared camera-spectrograph (Nadjip et al., 2017) mounted on the 2.5-m telescope. A detailed description of the observational procedure and data reduction can be found in Tatarnikov et al. (2023). The following stars from the 2MASS catalogue were used as photometric standards: 2MASS J21222481+4927597, 2MASS J21220594+4925335, and 2MASS J21222134+4925561. Their magnitudes were transformed into the MKO-NIR system using the color-term equations provided in Tatarnikov and Tatarnikov (2023). Additionally, we observed the source in the m) and m) bands – close to the corresponding MKO-NIR bands – using the LMP camera (Zheltoukhov et al., 2024). Photometric reference stars from the list of Shenavrin et al. (2011), positioned at the same airmass as the target during the observations, were employed as comparison stars. The results of our NIR photometry are presented in Table A2.
Polarimetric observations in the bands were carried out with the 2.5-m telescope using the speckle polarimeter (Safonov et al., 2017). The instrument is a dual-beam polarimeter equipped with a rotating half-wave plate as the modulator and a polarizing beamsplitting cube as the beamsplitter. The primary detector is a fast, low-noise CMOS camera (Hamamatsu ORCA-Quest). For each photometric band, we obtained a series of 4000 frames, each with an exposure time of 60 ms. During the acquisition, the half-wave plate rotated continuously at an angular velocity of 60∘ s-1.
The acquired images were corrected for bias offset and demodulated using the known position angle of the half-wave plate for each frame. Aperture photometry was then performed on the resulting demodulated series, taking into account the possible contribution from polarized background emission. The resulting flux measurements were used to estimate the Stokes parameters in the instrumental reference frame. Transformation to the equatorial reference frame and correction for instrumental polarization of the telescope (observations were carried out in the Nasmyth-2 focus) were performed following the procedure described in Safonov et al. (2017). The results are presented in Table A3 and Fig. 4.
Low-resolution spectra of IRAS 21204+4913 were obtained with the Transient Double-beam Spectrograph (TDS). A detailed description of the instrument and the data reduction procedure can be found in Potanin et al. (2020). Here we only note that the spectral resolving power of the TDS is in the red channel ( m) and in the blue channel ( m) (Belinski et al., 2023).
High-resolution spectra (HRS, over the wavelength range Å) were obtained with the ‘‘Raduga’’ spectrograph – an échelle fiber-fed spectrograph with a fiber diameter of . Data reduction was performed in the standard manner. Wavelength calibration was carried out using a hollow-cathode lamp (HCL). Spectral extraction followed the method described in Bolton and Schlegel (2010), employing HCL emission lines to construct a two-dimensional instrumental line-spread function. The high spatial resolution of the detector ( pixels) enables sufficiently detailed modeling of the line profile to solve the inverse problem of spectrum extraction. Wavelength calibration frames were acquired immediately before and after each science exposure, and the dispersion solution was interpolated to the mid-exposure time of the science frame. Stability tests using radial velocity standard stars (HIP 43726 and HIP 17378) showed that the local calibration accuracy is approximately km s-1 – this represents the scatter of individual spectral lines in the Raduga spectra relative to a reference spectrum obtained with the ESPRESSO spectrograph (Pepe et al., 2021) at . The random scatter in the mean radial velocities of the standard stars observed on different nights is km s-1.
All wavelengths and observation times for both spectrographs have been corrected to the barycenter of the Solar System. The log of spectral observations is presented in Table 1.
| date | object | BJD | exposure | PA |
| 246… | sec. | ∘ | ||
| HRS | ||||
| 2025-12-04 | 0 | 1014.23 | 3600 | - |
| 2025-12-06 | 0 | 1016.15 | 3600 | - |
| 2025-12-08 | 0 | 1018.28 | 3600 | - |
| 2025-12-19 | 0 | 1029.25 | 7200 | - |
| 2026-01-02 | 0 | 1043.21 | 7200 | - |
| TDS | ||||
| 2025-12-04 | 0 | 1014.26 | 600 | 70.8 |
| 2025-12-06 | 0 | 1016.18 | 600 | 85.7 |
| 2025-12-18 | 0 | 1028.19 | 1800 | |
| 2025-12-18 | 0 | 1028.17 | 1800 | 148.0 |
| 2026-01-03 | 0 | 1044.19 | 600 | 67.5 |
| 2025-12-20 | 1 | 1030.25 | 600 | 91.5 |
| 2026-01-29 | 1 | 1070.19 | 2400 | 56.7 |
| 2025-12-20 | 2 | 1030.25 | 600 | 91.5 |
| 2025-12-18 | HH | 1028.23 | 900 | 158.0 |
The object designations follow those in Fig. 1. HH denotes Herbig-Haro objects.
On December 3, 2025, we also obtained images of the IRAS 21204+4913 surroundings using the 4K4K CCD camera mounted on the 2.5-m telescope. Observations were carried out in the H filter ( nm, nm, a total exposure time of min), and in the adjacent continuum filter Hbc ( nm, nm, min).
3 Results
3.1 Image of the surroundings of IRAS 21204+4913 in the H line
Fig. 1 shows the difference image of the IRAS 21204+4913 surroundings obtained in the Halp and Halpbc filters. IRAS 21204+4913 is marked in the figure as object 0 at the southeastern edge of the dark cloud D 2944. On the southern and southwestern edges of the same cloud, approximately 50′′ and 227 west of IRAS 21204+4913, lie two stars that appear as bright spots, indicating strong H emission in their spectra (see also Barentsen et al. 2014). These stars – Gaia DR3 2171042269468334080 and Gaia DR3 2171089101807867520 – are labeled in Fig. 1 as objects 1 and 2, respectively. Hereafter, we refer to them as stars S1 and S2.
As can be seen from Fig. 1, a group of emission nebulae adjoins the northern part of the dark cloud D 2944. These nebulae lie in the direction from IRAS 21204+4913 at the projected distance of According to the spectrum shown in Fig. 2, the brighter, irregularly shaped nebulae – presumably Herbig-Haro (HH) objects – are superimposed on an extended emission background. The spectra of the background gas and the HH objects differ in two key respects. First, the [N ii] doublet lines are present in the spectrum of the background gas but absent in that of the HH objects. Second, the barycentric radial velocity of the background gas is km s-1, whereas that of the brighter clumps is km s-1. In both the background and HH spectra, the [S ii] Å line is significantly stronger than the Å line, implying an electron density cm-3 in the line-emitting regions (Proxauf et al., 2014).
3.2 Photometry
Fig. 3 shows the evolution of the brightness of IRAS 21204+4913 in the optical and NIR bands during our observation period, along with several of its color indices. It is evident that the rapid brightening in the optical (by nearly during November 2025; Stecklum 2025) slowed down significantly in December and essentially ceased by January. Throughout our monitoring period, the source’s brightness variations from 0.37 to 4.8 m occurred nearly simultaneously. More precisely, after December 10, 2025, the color indices and remained constant within the photometric uncertainties, while the color initially decreased slightly and then stabilized.
3.3 Polarimetry
It follows from Table A3 and Fig. 4 that the emission from IRAS 21204+4913 is strongly polarized in the bands. At each epoch, the degree of polarization increases with wavelength , while the polarization angle is constant within the measurement uncertainties. During the course of our observations, the degree of polarization in the bands increased, reaching in the band. The polarization angle also appears to have increased slightly during the observation period. This conclusion is quantitatively supported by a weighted least-squares fit (weights ) to the time dependence , which yields a slope that differs from zero at the level of .
3.4 Spectra
Fig. 5 shows low-resolution spectra of IRAS 21204+4913 obtained with TDS on different dates. Overall, the spectrum of the star resembles those of FUors (Fischer et al., 2023): it is dominated by absorption features, including the Li i 6708 Å line, and exhibits both high-excitation atomic lines (e.g. He i, O i, Si ii) and molecular TiO bands.
Another distinctive feature of the IRAS 21204+4913 spectrum that makes it resemble those of FUors (Herbig, 1975, Fischer et al., 2023) is the presence, in the blue wing of certain strong lines – particularly H (see Fig. 6) – of an extended absorption component. This fact indicates that the star, like FUors, has a gas shell that expands at velocities of km s-1, i.e., a wind.
Nevertheless, the spectrum of IRAS 21204+4913 exhibits features that are atypical for FUors. First, as seen in Fig. 5, the TiO molecular bands appear in emission rather than absorption. Second, there is a clear dependence both of the line width and radial velocity on the excitation energy (Fig. 7) – a behavior not previously observed in classical FUors, to the best of our knowledge.
The IRAS 21204+4913 spectrum exhibits several emission features. Firstly, as can be seen in Fig. 7, an emission component is present in the H line profile in all the spectra obtained by us. Furthermore, two emission lines are visible in the stellar spectrum: [S ii] 6731 Å (see Fig. 5) and [Ca ii] 7291 Å. However, both are very weak, and the profile of the latter, as well as that of the [S ii] 6716 Å line, is distorted by a neighboring absorption feature. Therefore, we refrain from analyzing the detailed shapes of these lines, although we can state that the radial velocities of their centroids is We also note that the [O i] 5577, 6300, and 6363 Å lines present in the IRAS 21204+4913 spectrum are of telluric origin. Another emission feature, observed in all spectra of the star near Å (Fig. 8), probably corresponds to the fluorescent Fe ii] 6516.08 Å line.444Transition , whose upper level is pumped by radiation from the levels of the ground configuration of Fe ii (Kramida et al., 2024). To the best of our knowledge, this line has not been previously reported in the spectra of FUors, although it is present in the spectra of some CTTS, e.g., RW Aur (Dodin et al., 2012).
3.5 Radial velocity estimation
In the previous section we showed that the radial velocity of IRAS 21204+4913 varies with the excitation potential of the spectral lines. To circumvent this complication and determine the stellar radial velocity , we proceeded as follows.
There are at least two CTTS in the vicinity of IRAS 21204+4913, labeled as objects 1 and 2 in Fig. 1. As shown in Table 1, we obtained two spectra of the star S1 and one spectrum of the star S2. The spectra of both stars correspond to a late M spectral type, exhibit significant veiling, and display strong emission lines of hydrogen, Ca ii H, K, and He i. [O i] emission line is also present in the spectrum of the S1 star (Fig. 9). For star S1, the equivalent width (EW) of H line varied from 100 Å (December 20, 2025) to 55 Å (January 29, 2026). In the spectrum of the S2 star, the H line exhibits a double-peaked profile with Å.
We have at our disposal a TDS spectrum of FN Tau – a CTTS of similar spectral type – which we used as a template for determining of S1 and S2 stars. The radial velocity was measured in the wavelength range Å, where the most prominent absorption features – TiO molecular bands – are located. The observed spectra of S1 and S2 were fitted to the FN Tau template using a least-squares minimization with three free parameters: radial velocity, veiling, and a scaling factor to match the mean flux levels. After correcting for the barycentric radial velocity of FN Tau, km s-1 (Jönsson et al., 2020), we obtained the following barycentric radial velocities: for S1: km s-1 (December 20, 2025) and km s-1 (January 29, 2026); for S2: km s-1 (December 20, 2025).
According to the Gaia DR3 catalogue (Gaia Collaboration et al., 2021), both stars lie at the same distance from Earth ( pc) and have similar proper motions within the uncertainties.555The catalogue does not provide distance or proper motion data for IRAS 21204+4913. Based on these data, we conclude that S1, S2, and IRAS 21204+4913 are members of a compact star-forming region associated with the dark cloud D 2944. In such a scenario, the radial velocities of the stars in this region should be similar. Therefore, our estimate of km s-1 for IRAS 21204+4913 appears reasonable.
3.6 Historical light curve
To investigate the long-term variability of IRAS 21204+4913, we estimated its brightness from archival photographic plates in the SAI MSU collection (see Section 2) and constructed a century-long light curve, shown in Fig. 10.
The star was first reliably detected on a photographic plate obtained on May 10, 1899 at a magnitude of It remained undetected thereafter until October 4, 1948, when it brightened to Because IRAS 21204+4913 lies close to a dark nebula, suitable comparison stars were relatively distant, which reduced the photometric accuracy. On the night of October 6-7, 1948, eleven photographic exposures of the field were obtained over a span of six hours, allowing us to assess the precision of our visual magnitude estimates. The average uncertainty is likely
As can be seen in Fig. 10, the star’s brightness increased from October 1948 until mid-August 1949, after which it began to decline. The last detection of the star on the photographic plates occurred on November 17, 1952.
4 Discussion
4.1 IRAS 21204+4913 and HH objects
The high degree of polarization of IRAS 21204+4913 and the wavelength independence of the polarization angle suggest that the polarization arises from scattering by dust. On p. 4.4 we will present evidence that this dust resides in an outflowing shell, that is a dusty wind. If the wind is axially symmetric, the polarization angle should be oriented perpendicular to the wind’s symmetry axis (Dodin et al., 2019). Therefore, from Fig. 4 it follows that the position angle of the wind axis is or
The second value is in excellent agreement with the direction from IRAS 21204+4913 to the HH objects: (see Fig. 1). We regard this alignment as strong evidence that the detected HH objects constitute part of a jet driven by the star under study. The barycentric radial velocity of the star is positive, whereas the HH objects exhibit negative radial velocities – see Figs. 7 and 2, respectively. This difference implies that the HH objects are moving toward Earth while simultaneously receding from the star.
4.2 Extinction toward IRAS 21204+4913
The absorption spectrum of IRAS 21204+4913 resembles those of A–F giants and supergiants, for example, V399 Car (A9 I) and Sco (F1 II). This suggests that the spectral energy distribution (SED) in the optical band should also be consistent with stars of these spectral types. We compared the observed SED of IRAS 21204+4913 with the SED of A–F giants and supergiants from the catalog of Pickles (1998) and found that, under the standard interstellar extinction law (; Cardelli et al. 1989), the extinction cannot exceed (the value is achieved for an A0 I star). For the F0 II spectrum – the closest available match to A9 I and F1 II in the catalog – we obtain
For larger values of , a kink appears near 4700 Å, which is not observed in the spectrum of IRAS 21204+4913 (Fig. 5). The kink becomes even more pronounced when hotter stars are used as spectral templates, ruling out the possibility of stronger reddening combined with steeper dereddened continuum for IRAS 21204+4913 than that of A–F giants. On the contrary, the presence of TiO emission bands indicates a contribution from radiation significantly cooler than that of A–F stars, which flattens the intrinsic SED by enhancing its red portion. Accounting for this cool component would lead to a lower estimate of
In the spectrum of IRAS 21204+4913 we found several diffuse interstellar bands (DIBs) from the list compiled by Herbig (1995). It is reasonable to assume that these bands arise in the portion of the dark nebula D 2944 that partially obscures the star. In particular, the EWs of the 5797 and 6714 Å DIBs are measured to be and Å, respectively, with barycentric radial velocities of and km s-1, respectively. Using the empirical relations between the EWs of these DIBs and extinction (Lan et al., 2015), we derive in the range , accounting for the observed scatter in the correlations, and when using the mean calibration. We also note the absence of detectable interstellar absorption in the Na i D and K i 7699 Å lines.
On the other hand, Stecklum (2025) reports that the Swift observatory did not detect X-ray emission from IRAS 21204+4913 in the keV band on November 27 and December 10, 2025, and on this basis concludes that However, during our observation period the optical spectrum of the star was nearly flat (Figs. 3 and 5). If such a high extinction were applied, the dereddened luminosity in the m range alone would reach L⊙, which is unrealistically large. This strongly suggests that the estimate by Stecklum (2025) applies only to the region where the X-ray emission originates, which is obscured by the intense dusty wind as in the case of FU Ori (Skinner et al., 2010).
Thus, we adopt that the extinction in the direction of IRAS 21204+4913 which is consistent with the estimate for D 2944 nebula as a whole (Dobashi, 2011). For such extinction, the maximum expected band interstellar polarization would be about 5.5 % (Hiltner, 1956), which is much less than the observed value. Besides, interstellar polarization typically exhibits a maximum near m. This gives reason to believe that the polarization of the object’s radiation occurs not in the interstellar medium, but in the dust shell of the star.
According to the catalog of Paegert et al. (2021), the band magnitude of the star prior to the 2025 outburst was 666Thus, the amplitude of the outburst in the band exceeded . Taking into account the adopted value of and the distance to the star of 500 pc, this corresponds to an absolute magnitude of Since eruptive YSOs are known to be younger than 10 Myr (Fischer et al., 2023), evolutionary models from Baraffe et al. (2015) imply that only stars with masses below 0.5 M⊙ can have absolute magnitudes fainter than this value of
4.3 Variability of the H line profile
We did not observe any significant differences in the IRAS 21204+4913 spectra obtained on different dates. The only exception is the profiles of the Balmer H i lines, especially the H line. We already noted (Sect. 3.4) that in all our spectra, starting from December 4, the H line has an emission component, which, however, is absent in the spectra obtained by Kochkina et al. (2025) on November 25 and 27. As can be seen from Fig. 6, the H line displays a P Cyg profile, in which the intensity of the redshifted peak increases with time, and the emission component in the blue absorption wing near km s-1 also strengthens. Note that Errico et al. (2003) also observed the H profile variations in FU Ori on timescales of several days, and Powell et al. (2012) even concluded that these changes appear to be periodic.
4.4 The ‘‘redshift’’ effect of spectral lines
It is reliably established that the spectrum of a FUor is formed in the protoplanetary disk, whose accretion luminosity greatly exceeds the luminosity of the central star (Fischer et al., 2023). Within this framework, the presence of lines with very different excitation potentials in the FUor spectra is explained by the radial decrease of the effective temperature of the disk with increasing distance from the rotation axis . Since the rotational velocity also decreases with , the model predicts that the observed line width – e.g. the full width at half maximum (FWHM) – should be smaller for lines with lower (Hartmann and Kenyon, 1996).
Meanwhile, Herbig (1989) already noted that, for all FUors known at the time, no dependence of on excitation energy was observed, at least in the visible range – see also Herbig et al. (2003), Petrov and Herbig (2008). The lack of correlation between and may indicate that the regions of the disk responsible for optical lines formation do not follow Keplerian rotation and/or that the FUor wind significantly distorts the observed line profiles (Petrov and Herbig, 2008).
We will not discuss these reasons in further detail, because in the case of IRAS 21204+4913 a clear correlation between and is indeed present. Moreover, as shown in Fig. 7, the radial velocity of the lines also increases with , a behavior not predicted by standard disk accretion models.
It can be seen from Fig. 7 that both dependencies arise because the red wing of the line becomes increasingly extended with higher . We believe that the change in line profiles is associated with the scattering of radiation (Grinin et al., 2012), a process whose significant role in shaping the spectrum of IRAS 21204+4913 is strongly supported by the high degree of polarization observed.
To illustrate this effect, consider a dusty wind in which the material moves radially outward at a constant velocity . A photon emitted by the central star at frequency will be Doppler-shifted in the rest frame of a dust grain to where is the speed of light. After a single coherent scattering event, the photon reaches the observer with frequency where is the angle between the radius vector pointing from the dust grain to the star and the line of sight. Eventually, the original wavelength is shifted by an amount
toward longer wavelengths. The effects of line shift and broadening due to multiple scattering are discussed in detail by Grinin et al. (2006).
The dependence of the line width on the excitation potential can be associated not only with the thermal structure of the accretion disk, but also with differences in the scattering parameters for disk regions with different temperatures. Different scattering geometries and/or scattering optical depth lead to different distortions of spectral lines. In other words, the observed spectrum depends on the geometry and kinematics of both the emitting region and the dusty wind, as well as on dust properties (Grinin et al., 2012).
Having no information on these parameters, as well as spectropolarimetric data for IRAS 21204+4913, we are forced to limit ourselves to a qualitative interpretation of the ‘‘redshift’’ effect.
4.5 Accretion rate estimate
The phenomenological division of eruptive YSOs into FUors and EXors, described in the Introduction, reflects fundamental differences in the accretion regimes of the protoplanetary disk’s matter onto the central star (Hartmann et al., 2016). EXors exhibit so-called magnetospheric accretion, similar to that in CTTSs, but with a much higher accretion rate For accretion rates typical of FUors ( M⊙yr-1), the protoplanetary disk “pushes through” the stellar magnetic field and makes direct contact with the stellar surface. Consequently, energy is released in a boundary layer near the star rather than in an accretion shock, as is the case for EXors and CTTSs.
Fig. 12 shows the SED of IRAS 21204+4913 in the m range, uncorrected for extinction. The SED is constructed from photometric data obtained at the beginning (December 9, 2025) and the end (January 29, 2026) of our observations (see Tables A1 and A2). During this period, the total flux in the considered wavelength range increased by approximately %, reaching a value of erg s-1cm-2. At a distance of 500 pc, this corresponds to a luminosity of L
For comparison, Fig. 12 also shows the pre-outburst SED of the star, constructed from archival data obtained at different epochs from the 2MASS (Skrutskie et al., 2006), WISE (Furlan et al., 2011), AKARI (Ishihara et al., 2010), and IRAS (Neugebauer et al., 1984) surveys. From these data we find that the pre-outburst bolometric luminosity of the star was L with the range m accounting for only %. It is unlikely that the flux at wavelengths m remained unchanged during the outburst. Therefore, it is reasonable to assume that the bolometric luminosity during this period was at least twice the value inferred from the m flux alone, reaching L
If this luminosity is powered by accretion onto a star of mass and radius then from the relation (Shakura and Sunyaev, 1973) we obtain that the accretion rate at that time must have exceeded
According to Baraffe et al. (2015), for young stars younger than 10 Myr with masses M⊙ (see p. 4.2), the ratio Therefore, we conclude that during our observation period the accretion rate was M⊙ yr-1.
4.6 Activity of the star before the 2025 outburst
Kochkina et al. (2025) found no evidence in the ASAS-SN data for strong outbursts of IRAS 21204+4913 between 2015 and October 2025. However, Stecklum (2025), relying on data from the Gaia (Mowlavi et al., 2021) and NEOWISE (Mei et al., 2023) observations, reports that before the outburst, IRAS 21204+4913 experienced long-term brightness variations of small amplitude in both the visible (after 2015) and NIR (after 2010) spectral regions. Stecklum (2025) also noted that a strong H line was observed in the spectrum of the star before 2025 (Barentsen et al., 2014) and suggested that intense accretion of circumstellar matter was already occurring before the 2025 outburst. This is further supported by the shape of the SED (Fig. 12) and the high luminosity of the star during that period.
Finally, we note (Section 3.6) that IRAS 21204+4913 underwent a prolonged brightness increase in the past. As can be seen from Fig. 10, the star experienced an outburst in 1948, during which its brightness rose to and then declined over approximately three years to a level of If this level represents the pre-outburst state of the star, then the amplitude of the 1948 outburst was
It’s difficult to say how similar the 1948 and 2025 outbursts are: we do not know what the star’s spectrum was during the first outburst, how quickly its brightness increased in 1948, and how long it will continue to decline after the second outburst. Already now, IRAS 21204+4913 is approximately brighter in the band than it was in 1949, however, we cannot rule out that the true maximum of the 1948 outburst was missed — see Fig. 10. Nevertheless, the very fact that an outburst of comparable amplitude occurred in an object that currently exhibits many characteristic features of a FUor is highly significant.
Herbig (1977), based on statistical grounds, concluded that the FUor phenomenon in young stars should be recurrent – see also Hillenbrand and Findeisen (2015). Meanwhile, among eruptive YSOs, considered ‘‘bonafide’’ FUors in the Contreras Peña,Carlos et al. (2025) catalog, no confirmed cases of repeated outbursts have been observed to date. On the other hand, in a number of objects with some characteristic FUor-like features – for example, V1647 Ori (Aspin et al., 2009) or V899 Mon (Ninan et al., 2015) – the brightness dropped sharply a few years after the maximum, and then began to increase again. Unlike these objects, IRAS 21204+4913 had a much longer interval between outbursts – at least 77 years.
Not all mechanisms invoked to explain FUor outbursts (see Introduction) can account for this kind of phenomenon. However, some models explicitly predict the presence of ‘‘double’’ outbursts in FUors, e.g., Grigoryev and Demidova (2025). In this regard, we recall that in EXors, repeated prolonged outbursts are observed quite frequently – see, for example, Herbig (1977), Dodin et al. (2016), Giannini et al. (2015).
5 Conclusion
There is no doubt that IRAS 21204+4913 is an eruptive YSO in the transitional stage from a protostar to a young star. The object is embedded within the dark nebula D 2944, which is a compact star-forming region located at a distance of approximately 500 pc and hosting low-mass YSOs M. We estimate the extinction toward IRAS 21204+4913 to be
The outburst of IRAS 21204+4913, which began in late October 2025, resembles FUor-type eruptions – phenomena driven by a dramatic increase in the accretion rate of material from the protoplanetary disk onto the forming star – in several aspects:
A large outburst amplitude corresponding to an accretion rate of M⊙yr-1;
An absorption spectrum in the range m resembling those of A–F giants and supergiants, combined with the presence of TiO molecular bands;
An outflowing shell expanding at velocities up to km s-1, in which the H line with a P Cyg profile is formed;
A group of HH objects in the vicinity of IRAS 21204+4913, receding from it with radial velocities of km s-1.
At the same time, we have identified several distinctive features in IRAS 21204+4913. As the star brightened, its radiation became increasingly polarized, reaching % in the band by the end of our observations. We attribute this to the presence of a dusty wind. The scattering of the central source’s radiation by dust grains produces a dependence of both the FWHM and of spectral lines on their excitation potential – an effect not previously observed in FUors. Another notable feature of IRAS 21204+4913 is an outburst of comparable amplitude that occurred 77 years ago.
We also discovered for the first time an emission line at Å and TiO molecular emission bands in the spectrum of an eruptive YSO with an absorption-dominated continuum. It is difficult to assess how critical these emission features are for understanding the outburst mechanism of IRAS 21204+4913. However, we note that the CO bands near m – typically seen in absorption in classical FUors – are observed in emission in the eruptive YSOs V1647 Ori and IRAS 06297+1021W (Connelley and Reipurth, 2018), which led Contreras Peña,Carlos et al. (2025) to classify them as intermediate objects between FUors and EXors.
It is still premature to classify IRAS 21204+4913 as a ‘‘bonafide’’ FUor, since – by definition – this group includes only objects whose brightness fades significantly over timescales of years to decades after maximum light (Samus’ et al., 2017). However, the issue is not merely one of formal classification of eruptive YSOs, but rather of understanding the physical mechanisms responsible for the diversity of their observed properties. The nontrivial features exhibited by IRAS 21204+4913 provide strong grounds to assert that this young star warrants close attention. It would be highly desirable not only to repeat the observations presented here, but also to obtain spectropolarimetric data and to monitor the evolution of its SED at wavelengths beyond 5 m.
Acknowledgements
We thank the staff of the CMO SAI MSU for their assistance during the observations. We are also grateful to M. R. Gilfanov, D. A. Lashin, and A. N. Tarasenkov for valuable discussions. We gratefully acknowledge the use of data from the following databases in this work: SIMBAD (CDS, Strasbourg, France), the Astrophysics Data System (NASA, USA), the NIST Atomic Spectra Database (NIST, USA), and the Atomic Line List (University of Kentucky, USA).
The study was conducted under the state assignment of Lomonosov Moscow State University.
Data availability
The results of the photographic magnitude measurements are available upon request.
Appendix
| Date | UT | ||||||||||
| 05.12.2025 | 17:21 | 12.51 | 0.22 | 12.56 | 0.02 | 11.95 | 0.02 | 11.47 | 0.02 | 10.77 | 0.01 |
| 06.12.2025 | 16:46 | 12.47 | 0.22 | 12.52 | 0.02 | 11.90 | 0.02 | 11.41 | 0.02 | 10.72 | 0.01 |
| 08.12.2025 | 17:45 | 12.40 | 0.22 | 12.46 | 0.02 | 11.83 | 0.02 | 11.34 | 0.02 | 10.65 | 0.01 |
| 09.12.2025 | 15:27 | 12.31 | 0.22 | 12.40 | 0.02 | 11.78 | 0.02 | 11.31 | 0.02 | 10.62 | 0.01 |
| 10.12.2025 | 15:14 | 12.16 | 0.22 | 12.26 | 0.02 | 11.64 | 0.02 | 11.16 | 0.02 | 10.47 | 0.01 |
| 16.12.2025 | 16:00 | 11.90 | 0.22 | 12.00 | 0.02 | 11.36 | 0.02 | 10.88 | 0.02 | 10.21 | 0.01 |
| 18.12.2025 | 16:03 | 11.85 | 0.22 | 11.94 | 0.02 | 11.30 | 0.02 | 10.82 | 0.02 | 10.15 | 0.01 |
| 19.12.2025 | 16:34 | 11.82 | 0.22 | 11.92 | 0.02 | 11.28 | 0.02 | 10.80 | 0.02 | 10.14 | 0.01 |
| 20.12.2025 | 16:15 | 11.75 | 0.22 | 11.86 | 0.02 | 11.22 | 0.02 | 10.75 | 0.02 | 10.10 | 0.01 |
| 21.12.2025 | 15:42 | 11.69 | 0.22 | 11.82 | 0.02 | 11.20 | 0.02 | 10.72 | 0.02 | 10.08 | 0.01 |
| 23.12.2025 | 15:04 | 11.72 | 0.22 | 11.81 | 0.02 | 11.18 | 0.02 | 10.69 | 0.02 | 10.05 | 0.01 |
| 25.12.2025 | 16:40 | 11.66 | 0.22 | 11.75 | 0.02 | 11.11 | 0.02 | 10.63 | 0.02 | 10.00 | 0.01 |
| 02.01.2026 | 15:59 | 11.57 | 0.22 | 11.65 | 0.02 | 11.00 | 0.02 | 10.52 | 0.02 | 9.90 | 0.01 |
| 03.01.2026 | 14:53 | 11.54 | 0.22 | 11.64 | 0.02 | 11.01 | 0.02 | 10.53 | 0.02 | 9.90 | 0.01 |
| 04.01.2026 | 17:01 | 11.57 | 0.22 | 11.66 | 0.02 | 11.02 | 0.02 | 10.54 | 0.02 | 9.92 | 0.01 |
| 07.01.2026 | 14:55 | 11.60 | 0.22 | 11.69 | 0.02 | 11.04 | 0.02 | 10.57 | 0.02 | 9.94 | 0.01 |
| 12.01.2026 | 16:10 | 11.57 | 0.22 | 11.60 | 0.02 | 10.95 | 0.02 | 10.48 | 0.02 | 9.85 | 0.01 |
| 15.01.2026 | 16:49 | 11.60 | 0.22 | 11.64 | 0.02 | 11.01 | 0.02 | 10.51 | 0.02 | 9.90 | 0.01 |
| 20.01.2026 | 15:10 | 11.59 | 0.22 | 11.61 | 0.02 | 10.95 | 0.02 | 10.47 | 0.02 | 9.84 | 0.01 |
| 21.01.2026 | 15.16 | 11.58 | 0.22 | 11.62 | 0.02 | 10.97 | 0.02 | 10.49 | 0.02 | 9.88 | 0.01 |
| 28.01.2026 | 15:32 | 11.63 | 0.22 | 11.68 | 0.02 | 11.04 | 0.02 | 10.56 | 0.02 | 9.93 | 0.01 |
| 29.01.2026 | 15:24 | 11.63 | 0.22 | 11.69 | 0.02 | 11.03 | 0.02 | 10.55 | 0.02 | 9.95 | 0.01 |
| Date | UT | ||||||||||||
| 05.12.2025 | 17:10 | 3.48 | 0.02 | 2.87 | 0.02 | ||||||||
| 06.12.2025 | 14:40 | 3.46 | 0.02 | 2.77 | 0.05 | ||||||||
| 09.12.2025 | 19:05 | 9.2 | 0.1 | 8.10 | 0.07 | 6.62 | 0.04 | 5.32 | 0.04 | 3.35 | 0.05 | 2.67 | 0.04 |
| 25.12.2025 | 19:10 | 8.00 | 0.06 | 6.52 | 0.05 | 5.18 | 0.02 | 3.18 | 0.04 | 2.40 | 0.05 | ||
| 03.01.2026 | 14:40 | 3.19 | 0.02 | 2.37 | 0.03 | ||||||||
| 07.01.2026 | 14:20 | 5.04 | 0.10 | 3.28 | 0.04 | 2.55 | 0.06 | ||||||
| 29.01.2026 | 16:00 | 8.84 | 0.02 | 7.73 | 0.02 | 6.39 | 0.03 | 5.21 | 0.02 | 3.27 | 0.04 | 2.52 | 0.04 |
| Date | UT | ||||||||||||||||
| dd.mm.yyyy | hh:mm | % | % | ∘ | ∘ | % | % | ∘ | ∘ | % | % | ∘ | ∘ | ||||
| 04.12.2025 | 18:35 | 13.85 | 0.17 | 65.3 | 0.7 | ||||||||||||
| 09.12.2025 | 16:17 | 7.26 | 0.19 | 65.0 | 1.5 | 8.10 | 0.20 | 64.0 | 1.4 | 13.96 | 0.16 | 65.2 | 0.7 | ||||
| 18.12.2025 | 18:04 | 8.74 | 0.17 | 66.1 | 1.1 | 9.42 | 0.17 | 65.6 | 1.1 | 15.02 | 0.17 | 66.4 | 0.6 | ||||
| 21.12.2025 | 16:30 | 9.25 | 0.17 | 66.1 | 1.0 | 9.97 | 0.17 | 65.6 | 1.0 | 15.68 | 0.16 | 66.5 | 0.6 | ||||
| 03.01.2026 | 14:57 | 9.41 | 0.17 | 70.2 | 1.0 | 10.39 | 0.22 | 68.2 | 1.2 | 15.86 | 0.16 | 68.2 | 0.6 | ||||
| 07.01.2026 | 15:06 | 8.61 | 0.23 | 66.8 | 1.6 | 9.67 | 0.18 | 68.4 | 1.0 | 10.66 | 0.17 | 68.3 | 0.9 | 15.53 | 0.18 | 66.6 | 0.7 |